# Context-free graphs and their transition groups

### v2

@@ 4. Transition groups of inverse graphs and co-context-free groups
Let $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ and consider the map $\psi:G\to\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$ sending an element $g\in G$ represented by a word $u\in A^{*}$ to the corresponding element $(\pi_{1}(u),\ldots,\pi_{k}(u))$, where $\pi_{i}:A\to\mathcal{G}(\Gamma^{(i)})$ are the natural projection maps. Now by definition
$\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})=WP(G;A).$
Hence $\psi$ is a well defined homomorphism since an element $g=\mathds{1}$ in $G$ is equivalent to $u\in WP(G;A)$, so $u\in\mathcal{L}(\Gamma^{(i)})$ and thus $\pi_{i}(u)=\mathds{1}$ for all $i=1,\ldots,k$. Moreover, $\psi$ is injective since $\pi_{i}(u)=\mathds{1}$ for all $i=1,\ldots,k$ is equivalent to $u\in\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=WP(G;A)$. ∎
Before proving the next lemma we need to recall the following definition. For a language $L\subseteq A^{*}$, the shift of $L$, is the language $S_{A}(L)=\{xy\in A^{*}:yx\in L\}$ formed by all the words that may be obtained by a cyclic shift of some word of $L$. If it is clear from the context, we drop the subscript $A$ and we simply refer to $S(L)$. It is a well-known fact (see for instance Ex. 6.4 of [18]) that for a context-free language $L$, $S(L)$ is also context-free.
@@
Let $\mathcal{G}(\Gamma)$ be transition group of a complete context-free graph $\Gamma$ and let $g$ be a non trivial element of $\mathcal{G}(\Gamma)$. There exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$ and $v\cdot g\neq v$.
@@
Let $q\in V(\Gamma)$ be the vertex from which there is a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}p$ with $p\neq q$ that is not a circuit. Consider the infinite one-sided walk $\pi=q\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, which exists since $\Gamma$ is complete. In $\pi$ choose a vertex $v$ belonging to the frontier vertices $\Delta(v,x_{0})$ such that $\pi\subseteq\Gamma(v,x_{0})$ (it exists since every vertex belongs to some set of frontier vertices and $\pi\subseteq\Gamma=\Gamma(x_{0},x_{0})$). Equivalently, $v$ is chosen among the vertices $\pi$ whose distance from $x_{0}$ is minimum. Using the vertex $v$ we may find a cyclic-shift $yx$ of $xy=g$ such that $\pi$ factors as follows:
$\pi=q\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}p\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$(yx)$}}}}{{\relbar\joinrel\longrightarrow}}}z_{3}\ldots z_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$(yx)^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$
where $v=z_{k}$ for some $k\geq 1$. We claim that there is a walk $v=z_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$yx$}}}}{{\relbar\joinrel\longrightarrow}}}z_{k+1}$ in $\Gamma(v,x_{0})$ that is not a circuit. Indeed, if $z_{k}=z_{k+1}$, then by the determinism of $\Gamma$ we would have $q=p$, a contradiction. ∎
@@
Let $\Gamma$ be a context-free inverse $A$-digraph that is also complete, then $\mathcal{L}(\Gamma)\in{\bf CoCF}$.
@@
$N(q)=\{u\in A^{*}:q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ is a walk in }\Gamma_{i}\mbox{ with }p\neq q\}$
@@
Let us prove that the right hand side is contained in the left hand side. Note that every walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Gamma_{i}$ that does not label a circuit at $q$, has the same property in $\Gamma$ in any subgraph that is end-isomorphic to $\Gamma_{i}$, i.e., $u\in\mathcal{L}(\Gamma)^{c}$. Moreover, take the word $yx$ such that $xy=u\in N(q)$, and factor a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ with $p\neq q$ in $\Gamma$ as $q\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}p$. If $yx$ labeled a circuit at $v$, then by the determinism of $\Gamma$ we would have $p=q$, a contradiction. Hence, $yx\in S(N(q))$ does not label a circuit at $v$ in $\Gamma$, i.e., $yx\in\mathcal{L}(\Gamma)^{c}$.
Let us now prove the other inclusion. So let $g\in\mathcal{L}(\Gamma)^{c}$, by Lemma 4.5 there exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$ and $v\cdot g\neq v$. Hence, $g\in S(N(w))$ where $w$ is the frontier vertex in the end-cone type $\Gamma_{j}$ that is end-isomorphic to the end-cone $\Gamma(v,x_{0})$, and this concludes the proof of the other inclusion. To show that $\bigcup_{i=1}^{N}\left(\bigcup_{q\in\Delta_{i}}S\left(N(q)\right)\right)$ is context-free, it is enough to show that each $N(q)$ is context-free, since context-free languages are closed under the cyclic-shift operator, and a finite union of context-free languages is context-free. Now, for any $q\in\Delta_{i}$, $N(q)=L(\Gamma_{i},q)^{c}$ and since $\Gamma_{i}$ is, in particular, a context-free graph, we have that $L(\Gamma_{i},q)$ is a deterministic context-free language since by Proposition 3.5, $\Gamma_{i}$ may be seen as the configurations graph of an inverse pushdown automaton accepting $L(\Gamma_{i},q)$. Now, since deterministic context-free languages are closed under complementation, we conclude that $L(\Gamma_{i},q)^{c}$ is also context-free, and this concludes the proof of the proposition. ∎
@@
In the previous proof we have shown that each language $N(q)$ is a deterministic context-free language. If $S(N(q))$ were deterministic for all $q\in\Delta_{i}$, then, by taking the complement, we would deduce that $\mathcal{L}(\Gamma)$ is an intersection of finitely many context-free languages, i.e., $\mathcal{L}(\Gamma)$ is poly-context-free. However, it seems that the cyclic closure of a context-free language is generally far from being deterministic, making the class of transition groups of context-free inverse graphs quite different from that of poly-context-free groups. In fact, we are going to find an example of a transition group which is not poly-context-free (see Definition 10.4).
@@
Let $\Gamma$ be a context-free inverse $A$-digraph that is also complete. Then, the transition group $\mathcal{G}(\Gamma)$ is CoCF.
@@
If $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ are context-free inverse $A$-graphs that are complete, then, the transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is CoCF.
@@
In view of Theorem 4.9, we may define the subclass of CoCF groups formed by all the transition groups of a finite disjoint collection of context-free inverse $A$-digraphs, denoted by CF-TR.
Question 4.11.
@@
There is a natural characterization of such groups in terms of certain coverings. We say that a covering $\psi:\Lambda\to\Gamma$ is transition-preserving if for any for any walk $q^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ with $q^{\prime}\neq p^{\prime}$ in $\Lambda$ there is a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Lambda$ such that the corresponding walk $\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(p)$ in $\Gamma$ is not a circuit. We have the following fact.
@@
If $\psi:\Lambda\to\Gamma$ is a transition-preserving covering between two complete $A$-digraphs, then $\mathcal{G}(\Gamma)\simeq\mathcal{G}(\Lambda)$. In particular, a group $G$ is CF-TR if and only if there is a transition-preserving covering of the Cayley graph $\operatorname{Cay}(G;A)$ onto a context-free $A$-digraph.
@@
Consider the map $\phi:\mathcal{G}(\Lambda)\to\mathcal{G}(\Gamma)$ defined as follows: for any $g\in\mathcal{G}(\Lambda)$ and $u\in A^{*}$ which represents $g$, we define $\phi(g)=h$ where $h\in\mathcal{G}(\Gamma)$ is the element defined by $u$. It is a well-defined map since if $v$ represents $g$ too, then $uv^{-1}\in\mathcal{L}(\Lambda)$ labels a circuit everywhere in $\Lambda$, and so also in $\Gamma$ since $\psi$ is a covering. Hence, $v$ represents $h$ as well. It is routine to check that it is a morphism of groups which is also surjective. Let us prove that $\phi$ is injective. Let $u\in\mathcal{L}(\Gamma)$, by the path lifting property (Proposition 2.6), any circuit $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}v$ in $\Gamma$ lifts to a walk $v^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ with $\psi(v^{\prime})=\psi(v^{\prime\prime})=v$. Now, if $u\notin\mathcal{L}(\Lambda)$, by the transition-preserving property we would have that $u\notin\mathcal{L}(\Gamma)$, a contradiction.

### v3

@@ 4. Transition groups of graphs and co-context-free groups
Let $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$. Define the map $\psi:G\to\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$ by putting $g=[u]\mapsto(\pi_{1}(u),\ldots,\pi_{k}(u))$, where $\pi_{i}:A\to\mathcal{G}(\Gamma^{(i)})$ are the natural projections. By definition $WP(G;A)=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})$, so $\psi$ is well-defined. Moreover, $\psi$ is injective: if $\pi_{i}(u)=\mathds{1}$ for all $i$, then $u\in\bigcap_{i}\mathcal{L}(\Gamma^{(i)})=WP(G;A)$, hence $g=\mathds{1}$ in $G$. ∎
4.1. Transition groups of context-free graphs
We now show that the transition group of any context-free graph is CoCF.
Before proving the next lemma we need to recall the following definition. For a language $L\subseteq A^{*}$, the shift of $L$, is the language $S_{A}(L)=\{xy\in A^{*}:yx\in L\}$ formed by all the words that may be obtained by a cyclic shift of some word of $L$. If it is clear from the context, we drop the subscript $A$ and we simply refer to $S(L)$. It is a well-known fact (see for instance Ex. 6.4 of [19]) that for a context-free language $L$, $S(L)$ is also context-free. For the next lemma, we need the following definition. Let $\pi=v\xrightarrow{u^{\omega}}\ldots$ be an infinite walk, for some $u\in A^{*}$. We say that a vertex $p$ is colored by a prefix $u^{\prime}$ of $u$ if $u=u^{\prime}u^{\prime\prime}$ and
$\pi=v\xrightarrow{u^{s}}v^{\prime}\xrightarrow{u^{\prime}}p\xrightarrow{u^{\prime\prime}}p^{\prime}\xrightarrow{u^{\omega}}\ldots$
for some integer $s\geq 0$.
@@ 4.1. Transition groups of context-free graphs
Let $\mathcal{G}(\Gamma)$ be transition group of a complete context-free graph $\Gamma$ and let $g$ be a non trivial element of $\mathcal{G}(\Gamma)$. There exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $\pi=v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$\mathring{g}^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$. Moreover, for any vertex $p\in\pi$ colored by a prefix $g^{\prime}$ such that $g=g^{\prime}g^{\prime\prime}$ we have $p\cdot(g^{\prime\prime}g^{\prime})\neq p$.
@@
Since $g$ is nontrivial, there exists a vertex $q_{1}\in V(\Gamma)$ and a walk $q_{1}\xrightarrow{g}q_{2}$ with $q_{1}\neq q_{2}$. Consider the infinite one-sided walk $\pi^{\prime}=q_{1}\xrightarrow{g^{\omega}}\ldots$, which exists since $\Gamma$ is complete. Choose a vertex $v$ in $\pi^{\prime}$ belonging to the frontier vertices $\Delta(v,x_{0})$ such that $\pi^{\prime}\subseteq\Gamma(v,x_{0})$ (this is possible since every vertex belongs to some frontier set, and $\pi^{\prime}\subseteq\Gamma(x_{0},x_{0})=\Gamma$). Equivalently, $v$ is chosen among the vertices of $\pi^{\prime}$ that are at minimal distance from $x_{0}$. Using the vertex $v$, we may find a cyclic shift $\mathring{g}$ of $g$ such that $\pi=v\xrightarrow{\mathring{g}^{\omega}}\ldots\subseteq\Gamma(v,x_{0})$. Now, for any vertex $p\in\pi$ colored by a prefix $g^{\prime}$ of $g$, with $g=g^{\prime}g^{\prime\prime}$, it is straightforward to see that, by the determinism of $\Gamma$, the equality $p\cdot g^{\prime\prime}g^{\prime}=p$ would imply $q_{1}=q_{2}$, a contradiction. Hence, necessarily $p\cdot g^{\prime\prime}g^{\prime}\neq p$. ∎
@@
Let $\Gamma$ be a context-free $A$-graph that is also complete, then $\mathcal{L}(\Gamma)\in{\bf CoCF}$.
@@
$N(q)=\{u\in A^{*}:q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ is a walk in }\Gamma_{i}\mbox{ with }p\neq q\}$
@@
Let us prove that the right hand side is contained in the left hand side. Note that every walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Gamma_{i}$ that does not label a circuit at $q$, has the same property in $\Gamma$ in any subgraph that is end-isomorphic to $\Gamma_{i}$, i.e., $u\in\mathcal{L}(\Gamma)^{c}$. Moreover, take the word $yx$ such that $xy=u\in N(q)$, and factor a walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p$ with $p\neq q$ in $\Gamma$ as $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$x$}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y$}}}{{\relbar\joinrel\longrightarrow}}}p$. If $yx$ labeled a circuit at $v$, then by the determinism of $\Gamma$ we would have $p=q$, a contradiction. Hence, $yx\in S(N(q))$ does not label a circuit at $v$ in $\Gamma$, i.e., $yx\in\mathcal{L}(\Gamma)^{c}$.
Let us now prove the other inclusion. So let $g\in\mathcal{L}(\Gamma)^{c}$, by Lemma 4.7 there exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$\mathring{g}^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$ and $v\cdot\mathring{g}\neq v$. Hence, $g\in S(N(w))$ where $w$ is the frontier vertex in the end-cone type $\Gamma_{j}$ that is end-isomorphic to the end-cone $\Gamma(v,x_{0})$, and this concludes the proof of the other inclusion. To show that $\bigcup_{i=1}^{N}\left(\bigcup_{q\in\Delta_{i}}S\left(N(q)\right)\right)$ is context-free, it is enough to show that each $N(q)$ is context-free, since context-free languages are closed under the cyclic-shift operator, and a finite union of context-free languages is context-free. Now, for any $q\in\Delta_{i}$, $N(q)=L(\Gamma_{i},q)^{c}$ and since $\Gamma_{i}$ is, in particular, a context-free graph, we have that $L(\Gamma_{i},q)$ is a deterministic context-free language by Proposition 3.4. Now, since deterministic context-free languages are closed under complementation, we conclude that $L(\Gamma_{i},q)^{c}$ is also context-free, and this concludes the proof of the proposition. ∎
@@
In the previous proof we have shown that each language $N(q)$ is a deterministic context-free language. If $S(N(q))$ were deterministic for all $q\in\Delta_{i}$, then, by taking the complement, we would deduce that $\mathcal{L}(\Gamma)$ is an intersection of finitely many context-free languages, i.e., $\mathcal{L}(\Gamma)$ is poly-context-free. However, it seems that the cyclic closure of a context-free language is generally far from being deterministic, making the class of transition groups of context-free graphs quite different from that of poly-context-free groups. In fact, we are going to find an example of a transition group which is not poly-context-free (see Definition 10.4).
@@
Let $\Gamma$ be a context-free $A$-graph that is also complete. Then, the transition group $\mathcal{G}(\Gamma)$ is CoCF.
@@
If $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ are context-free $A$-graphs that are complete, then, the transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is CoCF.
@@
In view of Theorem 4.11, we may define the subclass of CoCF groups formed by all the transition groups of a finite disjoint collection of context-free $A$-graphs, denoted by CF-TR.
Question 4.13.
@@
There is a natural characterization of such groups in terms of certain coverings. We say that a covering $\psi:\Lambda\to\Gamma$ is transition-preserving if for any for any walk $q^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ with $q^{\prime}\neq p^{\prime}$ in $\Lambda$ there is a walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Lambda$ such that the corresponding walk $\psi(q)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}\psi(p)$ in $\Gamma$ is not a circuit. We have the following fact.
@@
If $\psi:\Lambda\to\Gamma$ is a transition-preserving covering between two complete $A$-graphs, then $\mathcal{G}(\Gamma)\simeq\mathcal{G}(\Lambda)$. In particular, a group $G$ is CF-TR if and only if there is a transition-preserving covering of the Cayley graph $\operatorname{Cay}(G;A)$ onto a context-free $A$-graph.
@@
Consider the map $\phi:\mathcal{G}(\Lambda)\to\mathcal{G}(\Gamma)$ defined as follows: for any $g\in\mathcal{G}(\Lambda)$ and $u\in A^{*}$ which represents $g$, we define $\phi(g)=h$ where $h\in\mathcal{G}(\Gamma)$ is the element defined by $u$. It is a well-defined map since if $v$ represents $g$ too, then $uv^{-1}\in\mathcal{L}(\Lambda)$ labels a circuit everywhere in $\Lambda$, and so also in $\Gamma$ since $\psi$ is a covering. Hence, $v$ represents $h$ as well. It is routine to check that it is a morphism of groups which is also surjective. Let us prove that $\phi$ is injective. Let $u\in\mathcal{L}(\Gamma)$, by the path lifting property (Proposition 2.6), any circuit $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}v$ in $\Gamma$ lifts to a walk $v^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ with $\psi(v^{\prime})=\psi(v^{\prime\prime})=v$. Now, if $u\notin\mathcal{L}(\Lambda)$, by the transition-preserving property we would have that $u\notin\mathcal{L}(\Gamma)$, a contradiction.